An electron travels at a constant speed toward the bottom of this page. It then enters a 3.50 × 105 T uniform magnetic field and experiences a maximum force of 5.15 × 10-8 N that points toward the right of the page.

a) What is the speed of the electron?
Ans: ___________ m/s

. b) What is the direction of the magnetic field?

(option a) into this page

(option b) toward the right of this page.

(option c) out of this page

(option d) toward the bottom of this page.

(option e) toward the left of this page.

(option f) toward the top of this page.
Answer #1

F=q*v*B*sin(\theta)

=>v =F/(e*B) =5.15*10^-8N/(1.6022*10^-19*3.5*10^5) =9.18*10^5m/s

c) out of page